# Deep learning in a bilateral brain with hemispheric specialization

### 05 Mar 2024 v7

@@ 4.1.3 Framework and computational resources
4.2 Visualizing network operation
For visualizations, we focussed on one hemisphere architecture, ResNet-9. We used two types of visualizations to better understand the operation of the bilateral network.
In order to be informative, we selected key scenarios that are distinct from each other and highlight the contribution of different parts of the network. The scenarios are:
  • Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect.
  • Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect.
  • Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect.
  • Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct.
The definition of ‘correct’ for the bilateral-specialized network, is that it was successful for both fine and coarse class labels.
Gradient camera (Grad-Cam) visualization
To understand how the specialized bilateral network exploits features from both hemispheres, we visualized gradient flow with respect to target convolutional layers [42, 43, 44] using the Grad-Cam library. The gradients reveal which areas of the image contribute most to the prediction; the gradient heatmap highlights the region of focus. We applied the visualization to each of the specialized hemispheres individually, and compared it to the bilateral network.
For a single hemisphere, the gradients were averaged over the 2nd convolutional layer of each residual block (2 convolutional layers with skip connections, see Figure 2) during prediction. In the case of the bilateral network, the gradients were averaged over both hemispheres during simultaneous prediction of both coarse and fine labels (the network had two heads).
@@ 4.2 Visualizing network operation
We investigated how left and right features are exploited by the bilateral network by analyzing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads. We did this by measuring the similarity of features for images of the same label. They are expected to have similar features, so measuring the feature similarity at different parts of the network should be revealing.
@@ 5.1 Accuracy
Quantitative results are shown in tabular form in Tables 1 and 2 and with plots in Figures 3 and 4.
The relative performance of models was very consistent between ResNet and VGG architectures. In summary, the bilateral architecture with specialization provided a performance boost for both fine and coarse classes over an individual specialized hemisphere (unilateral-specialized). Moreover, it outperformed the baselines, except for the 2-model ensemble. The 2-model ensemble consists of unilateral models, which were more effective than unilateral that were specialized in most cases. The exception was the VGG architecture with coarse labels, where the unspecialized unilateral model was not better than the specialized version, and the resulting 2-model ensemble did not outperform the specialized bilateral network.
@@
Unilateral (specialized for fine) 6.58 67.81 $\pm$ 0.31
Unilateral (specialized for coarse) 6.62 75.87 $\pm$ 0.70
@@
Table 1: Accuracy of ResNet-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
@@
Unilateral (specialized for fine) 129.18 51.87 $\pm$ 0.58
Unilateral (specialized for coarse) 128.85 66.86 $\pm$ 0.41
@@
Table 2: Accuracy of VGG-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
Figure 3: Accuracy of ResNet-based architectures. For brevity, if not specified, the network is unspecialized. The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
Figure 4: Accuracy of VGG-11-based models. For brevity, if not specified, the network is unspecialized. The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
5.2 Visualizing network operation
Visualizations of Grad-Cam and similarity distributions are shown for the selected scenarios described in Section 4.2.
The Grad-Cam visualizations are shown in Figures 5 to 8. In general, the features are more local in the left hemisphere and more global in the right hemisphere. The heads blend different aspects of features from the left and right hemispheres for the task. The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
@@ 5.2 Visualizing network operation
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 5: Grad-Cam for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect. The bilateral network appears to adjust the area of focus even when both the hemispheres’ features are situated external to the bosc.
@@
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 6: Grad-Cam for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect. Possibly due to the fact that there are multiple labels with wheels, the left failed to differentiate the tank by its wheels alone, however the bilateral appears to overcome this, using the right hemisphere’s features.
@@
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 7: Grad-Cam for Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect. The left hemisphere identified more local features of the bicycle compared to the right. The network head appear to exploit the left representations to make an accurate prediction.
@@
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 8: Grad-Cam for Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct. Both hemispheres identified local and global features. The network heads appear to over-compensate and focus on the region outside of the sofa, which is ultimately unsuccessful.
The cosine similarity results are shown in Figure 9 and Figure10. In the concatenated distribution, there is a strong correlation between similarity in left and concatenated, and right and concatenated. In contrast, there is no obvious correlation in the bilateral distribution. The network heads learned a non-linear transformation of the feature space.
@@
(b) Bilateral-specialized network
Figure 9: Cosine similarity distribution for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect. Many of the pairs (same label) have dissimilar (values closer to 0) features in left and right, seen in the density of points close to the origin. The concatenated features are also dissimilar, but the bilateral network elevates the similarity of many of the points. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
@@
(b) Bilateral-specialized network
Figure 10: Cosine similarity distribution for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect. Some regions with points that have low similarity in the left (at approximately x, y = 0.2, 0.2), have elevated similarity in the bilateral output, compared to concatenated. Also, for points with high similarity in the left (approximately 0.35, 0.1), the similarity is lower in the bilateral output. In other words, the bilateral network was able to exploit features in the right hemisphere over the left. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
@@ 6 Discussion
6.1 Key findings
We conducted experiments to implement and study bilateralism in Artificial Neural Networks. The effects of bilateralism were studied by comparing against baselines that captured distinct characteristics of the proposed bilateral network. We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialized for fine classes and the right for coarse classes.
@@ 6.1 Key findings
We found that there is an advantage of two hemispheres over one, demonstrated by the fact that Bilateral-specialized outperformed individual hemispheres alone. Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialization is important, shown by the benefit of bilateral-specialized over bilateral.
@@
Surprisingly, an unspecialized hemisphere (Unilateral) performed as well or better than specialized hemispheres (Unilateral-specialized) in most cases. Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
@@
Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results. A key difference with Bilateral-specialized, is that the component models were more effective to begin with. Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector. Future work could explore this in more detail.
@@
Both the specialized bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
@@
The novel finding of the paper is that a bilateral architecture can exploit specialized hemispheres, and the remainder of the paper examines this principles.
6.2 Computational cost
@@
Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost. The performance boost between unilateral and specialized bilateral models may appear modest for the additional cost. However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential. This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
6.3 How does the bilateral architecture work?
@@
The Grad-Cam images reveal that the left hemisphere extracts more localized features than the right. Different learning objectives enable them to capture different aspects of the environment. Collectively the set of features is greater than one network with one objective.
Interestingly, even though the left is explicitly trained on fine class labels, the features that it extracts are helpful for coarse classes. The inverse is true of the right.
@@
With reference to the cosine similarity visualization, before the network heads, the similarity of pairs of images of the same class is correlated with left and right. The similarity is no longer correlated after the network heads, which must implement some sort of non-linear transformation. In many cases, when left or right produce ineffective features, shown by low similarity between these images of the same class, the heads are able to produce features that have increased similarity. The network heads learn to combine the features from left and right, to different degrees, to produce better predictions. In some cases, the bilateral network makes a correct classification, even if both hemispheres are individually wrong.
@@
In summary, specialization creates a higher diversity of features. The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
6.4 Relation to biology
@@
The central finding that in our model, left and right hemispheres extracted local and non-local features, used together in fine tasks, provides one plausible suggestion for the prevalence of bilateral brains in nature. A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
6.5 Why it’s important
The key findings are a small step towards understanding and taking advantage of the characteristics of biological brains, and show the potential of bilateralism to improve ML/AI. More broadly, it’s a biological principle that warrants further investigation.
The fact that bilateralism has a material effect on the network as a whole, suggests that bilateralism should be considered for other cognitive neuroscience models, where it is usually ignored. One potentially fruitful area is extending the standard Hippocampal/Neocortical model, CLS [47, 48], by introducing asymmetries into the bilateral Hippocampal architecture. Indeed, evidence exists of asymmetric contributions of left and right Hippocampi to memory [49, 50].
Bilateralism could also inspire new priors or inductive biases to improve ML models.
6.6 Future work
Given that conventional ensembling was effective, and in addition differential specialization provided benefits, it would be interesting to examine the advantage of each. For example, does one of these methods take you further on different or more difficult datasets? Furthermore, are there ways to combine aspects of both methods, such ensembling bilateral models – would this be more effective than simply increasing the size of the ensemble for the same total number of hemispheres. Finally, can we leverage the specialized features in more effective ways, such as using a combiner network as opposed to simple concatenation of features.
A promising direction to improve the model as well as our understanding of the neuroscience, is to further investigate and model the neurobiology. There are several avenues. For example, replicating recurrent connectivity, more complex biologically inspired interactions between the hemispheres, mimicking known substrate differences between the hemispheres such as topological differences, resource allocation (see Figure LABEL:fig:bilateral_figure) and experiments on inducing specialization without supervision.
This study focused on one aspect of bilateralism with an image classification task. There are also interesting lateral effects for motor control. For example, in most people, the right hand is better at trajectories, and the left hand is better at position/posture (dynamic dominance theory [51]). Developing an architecture and experiments to explore motor control is an interesting area for future work.
One theory is that the observed differences between the hemispheres emerge from a more fundamental specialization for novelty (right) and routine (left) [4, 5, 52]. These ideas could be explored with an agent that can act in a dynamic environment. The right hemisphere could be a generalist that can perform unfamiliar tasks as a beginner, while the left becomes an expert over time. The agent would be able to adopt new tasks, without being inept at them. Currently, the field of Continual RL does not focus on avoiding poor performance, but rather maximizing the best performance. In real-life scenarios however, an agent must avoid death and serious injury to themselves and those around them (also relevant for physical robots and virtual artificial agents).
As mentioned earlier in this section, bilateralism could be incorporated into existing cognitive computational models, and the benefits of bilateralism could be studied within more powerful ML architectures.
@@ 7 Conclusion
We built a bilateral architecture with left and right neural networks, inspired by our bilateral brains. The hemispheres were trained to specialize on fine and coarse classes. They extracted specialized features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes. The specialized representations had benefits above the explicit objective of their individual hemispheres. The results demonstrate that small procedural changes to training can achieve specialization, and that specialization can be complementary and beneficial for certain tasks. The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialization. Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures. Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
Acknowledgments
In an earlier Masters project, co-supervised by Levin Kuhlmann, Amir Mayan explored similar ideas, which formed a valuable background to planning this project. Thanks to Punarjay Chakravarty for helpful discussions.
Author contributions
GK conceived and supervised the project, CR designed and implemented the experiments, GK and CR authored the manuscript, DR advised and assisted with the manuscript and EG introduced the concept and provided advice. All authors read and approved the final manuscript.

### 15 Mar 2024 v8

@@ 4.1.3 Framework and computational resources
4.2 Visualising network operation
We focused on one backbone, ResNet-9, and used two types of visualisations to better understand the operation of the bilateral network. To be informative, we selected key scenarios that are distinct from each other and highlight the contribution of different parts of the network, described below. The definition of ‘correct’ for the bilateral-specialised network, is that it was successful for both fine and coarse class labels.
  • Scenario 1: Bilateral-specialised network is correct, left and right hemispheres are incorrect
  • Scenario 2: Bilateral-specialised network and right hemisphere are correct, left hemisphere is incorrect
  • Scenario 3: Bilateral-specialised network and left hemisphere are correct, right hemisphere is incorrect
  • Scenario 4: Bilateral-specialised network is incorrect, left and right hemispheres are correct
Gradient camera (Grad-Cam) visualisation
To understand how the specialised bilateral network exploits features from both hemispheres, we visualised gradient flow with respect to target convolutional layers [43, 42] using the Grad-Cam library [44]. The gradients reveal which areas of the image contribute most to the prediction; the gradient heatmap highlights the region of focus. We applied the visualisation to each of the specialised hemispheres individually, and compared it to the bilateral network.
For a single hemisphere, the gradients were averaged over the 2nd convolutional layer of each residual block (2 convolutional layers with skip connections) during prediction. In the case of the bilateral network, the gradients were averaged over both hemispheres during simultaneous prediction of both coarse and fine labels (the network had two heads).
@@ 4.2 Visualising network operation
We investigated how left and right features are exploited by the bilateral network by analysing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads. We did this by measuring the similarity of features for images of the same label. They are expected to have similar features, so measuring the feature similarity at different parts of the network should be revealing.
@@ 5.1 Accuracy
The relative performance of models was very consistent between ResNet and VGG backbones, Tables 1 and 2 and Figures 2 and 3. Accuracy given as mean $\pm$ 1 standard deviation (shown as error bars in the plots) of 5 seeds. In summary, Bilateral-specialised outperformed all the baselines, except for the 2-model ensemble. The 2-model ensemble consists of unilateral models, which were more effective than specialised unilateral models in most cases. The exception was the VGG backbone with coarse labels, where the unspecialised unilateral model was not better than the specialised version, and the resulting 2-model ensemble did not outperform Bilateral-specialised.
@@
Unilateral-specialised 6.58 67.81 $\pm$ 0.31
Unilateral-specialised 6.62 75.87 $\pm$ 0.70
@@
Table 1: Accuracy of ResNet-based models.
@@
Unilateral-specialised 129.18 51.87 $\pm$ 0.58
Unilateral-specialised 128.85 66.86 $\pm$ 0.41
@@
Table 2: Accuracy of VGG-based models
Figure 2: Accuracy of ResNet-based models. For brevity, if not specified, the network is unspecialised.
Figure 3: Accuracy of VGG-11-based models. For brevity, if not specified, the network is unspecialised.
5.2 Visualising network operation
The Grad-Cam visualisations are shown in Figures 4 to 7 for the scenarios described in Section 4.2. In general, the features appear to be more local in the left hemisphere and more global in the right. The heads blend different aspects of features from the left and right, for the task. The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
@@ 5.2 Visualising network operation
(b) Left
(c) Right
(d) Bilateral-sp
Figure 4: Grad-Cam for Scenario 1: The bilateral-specialised (abbreviated to bilateral-sp in the figure) network is correct, left and right are incorrect. The bilateral network appears to adjust the area of focus even when both the hemispheres’ features are situated external to the bosc.
@@
(b) Left
(c) Right
(d) Bilateral-sp
Figure 5: Grad-Cam for Scenario 2: The bilateral-specialised network and the right are correct, the left is incorrect. Possibly due to the fact that there are multiple labels with wheels, the left failed to differentiate the tank by its wheels alone, however the bilateral model appears to overcome this, using the right’s features.
@@
(b) Left
(c) Right
(d) Bilateral-sp
Figure 6: Grad-Cam for Scenario 3: The bilateral-specialised network and the left are correct, the right is incorrect. The left hemisphere identified more local features of the bicycle compared to the right. The network head appear to exploit the left representations to make an accurate prediction.
@@
(b) Left
(c) Right
(d) Bilateral-sp
Figure 7: Grad-Cam for Scenario 4: The bilateral-specialised network is incorrect, the left and right hemispheres are correct. Both hemispheres identified local and global features. The network heads appear to over-compensate and focus on the region outside of the sofa, which is ultimately unsuccessful.
The cosine similarity results are shown in Figure 8 and Figure9. In the concatenated distribution, there is a strong correlation between similarity in left and concatenated, and right and concatenated. In contrast, there is no obvious correlation in the bilateral distribution. The network heads appear to have learned a non-linear transformation of the feature space.
@@
(b) Bilateral-specialised network
Figure 8: Cosine similarity distribution for Scenario 1: The bilateral-specialised network is correct, left and right are incorrect. Many of the pairs (same label) have dissimilar (values closer to 0) features in left and right, seen in the density of points close to the origin. The concatenated features are also dissimilar, but the bilateral network elevates the similarity of many of the points. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
@@
(b) Bilateral-specialised network
Figure 9: Cosine similarity distribution for Scenario 2: The bilateral-specialised network and the right are correct, the left is incorrect. Some regions with points that have low similarity in the left (at approximately x, y = 0.2, 0.2), have elevated similarity in the bilateral output, compared to concatenated. Also, for points with high similarity in the left (approximately 0.35, 0.1), the similarity is lower in the bilateral output. In other words, the bilateral network was able to exploit features in the right hemisphere over the left. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
@@ 6 Discussion
We found that there is an advantage of two hemispheres over one, demonstrated by the fact that the specialised bilateral architecture (Bilateral-specialised) outperformed individual hemispheres (Unilateral-specialised and Unilateral). Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialisation is important, shown by the benefit of Bilateral-specialised over Bilateral.
@@
Surprisingly, unspecialised hemispheres (Unilateral) performed as well or better than specialised hemispheres (Unilateral-specialised) in most cases. Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
@@
Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results. A key difference with Bilateral-specialised, is that the component models were more effective to begin with. Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector. Future work could explore this in more detail.
@@
Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
@@
The novel finding of the paper is that a bilateral architecture can exploit specialised hemispheres, and the remainder of the paper examines this principle.
Computational cost
@@
Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost. The performance boost between unilateral and specialised bilateral models may appear modest for the additional cost. However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential. This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
How does the bilateral architecture work?
@@
The Grad-Cam images reveal that the left hemisphere extracts more localised features than the right. Different learning objectives enable them to capture different aspects of the environment. Collectively the set of features is greater than one network with one objective. Interestingly, even though the left is explicitly trained on fine labels, the features that it extracts are helpful for coarse classes. The inverse is true of the right.
@@
With reference to the cosine similarity visualisation, before the network heads, the similarity of pairs of images of the same class is correlated with left and right and the similarity is no longer correlated after the network heads; which suggests that the heads implement a non-linear transformation. In many cases, when left or right produce ineffective features, shown by low similarity between embeddings of these images of the same class, the heads are able to produce features that have increased similarity. The network heads learn to combine the features from left and right, to different degrees, to produce better predictions. In some cases, the bilateral network makes a correct classification, even if both hemispheres are individually wrong.
@@
In summary, specialisation creates a higher diversity of features. The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
Relation to biology
@@
The central finding, that left and right hemispheres extracted local and non-local features, to improve performance on a given task, provides one plausible suggestion (and testable hypothesis) for the prevalence of bilateral brains in nature. A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
@@ 7 Conclusion
Inspired by our bilateral brains, we built a bilateral architecture with left and right neural networks using two CNN backbones and compared them to baselines that captured distinct characteristics of the proposed bilateral network. We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialised for fine classes and the right for coarse classes. The hemispheres were trained to specialise on fine and coarse classes accordingly. They extracted specialised features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes. The specialised representations had benefits above the explicit objective of their individual hemispheres. The results demonstrate that small procedural changes to training can achieve specialisation, and that specialisation can be complementary and beneficial for certain tasks. The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialisation. Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures. Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
Acknowledgements
Thanks to Punarjay Chakravarty for helpful discussions.